A stock DEF has the following payoffs probabilities: Probability 0.3 0.45 0.25 Payoff $50 $60 $90...

Question 4 Saved

A stock DEF has the following payoffs probabilities:

Probability

0.3

0.45

0.25

Payoff

$50

$60

$90

What is the Variance of payoffs of the stock?

Note: The numbers for this question may have been changed from the previous question.

Give your answer for Variance as a number and not a percentage.

Your Answer:
Answer #1

Expected payoff=Respective payoff*Respective probability

=(0.3*50)+(0.45*60)+(0.25*90)=$64.50

probability Payoff probability*(Payoff-Mean)^2
0.3 50 0.3*(50-64.5)^2=63.075
0.45 60 0.45*(60-64.5)^2=9.1125
0.25 90 0.25*(90-64.5)^2=162.5625
Total=234.75

Standard deviation=[Total probability*(Payoff-Mean)^2/Total probability]^(1/2)

=15.32(Approx)

Hence variance=Standard deviation^2

=234.75